A silver necklace originaly priced at $45.50 is on sale for 1/5 off. what is the sale price?
i can't understand how to to calculate this. please show me in detail.
thanku!

Answer by quentin252(19)

ok,
the necklace priced at 45.50 is 1/5 off.
so 1/5 of 45.50 is 1/5 * 45.50 = 9.10
so this is the amount off the original price.
so the final price = 45.50 - 9.10 = 36.40
good luck!!